OpenStudy (ashy98):

A. Given the equation 4x + 7 = 3(2x − 5), solve for the variable. Explain each step and justify your process. B. Sam solved a similar equation below. Is Sam's solution correct? Explain why or why not. 5x − 1 = 2(x + 4) 5x − 1 = 2x + 8 7x − 1 = 8 7x = 7 x = 1

OpenStudy (tkhunny):

Well, do it? What's preventing you from taking the first step?

OpenStudy (ashy98):

Part A: 4x + 7 = 6x - 15 4x - 6x = 15 - 7 (Collect all like terms) -2x = 8 (divide both side by -2) x = -4

OpenStudy (anonymous):

Part A: x=11 4x+7=6x-15 [Distributive Property] 4x=6x-22 [Subtract 7 from both sides] -2x=-22 [Subtract 6x from both sides] x=11 [Divide Both sides by -2] Part B: Sam is incorrect Sam got the first step right but he messed up the second, he added 2x on both sides to get 7x on the left when in fact he should've subtracted 2x to give him 3x on both sides